OpenStudy (anonymous):

determine an equation in simplified form, for a function with the zeroes of 2+ square root of 5 / 2- square root of 5 and 0., and with its graph passing through the point (2,20)

OpenStudy (anonymous):

\[2-\sqrt{5}, 2+\sqrt{5}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you are looking for an equation with specific roots right?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

\[(x-(2-\sqrt{5}))(x-(2+\sqrt{5}))\]

OpenStudy (anonymous):

there it is if you sub either of those numbers into it you will set the equation to zero

OpenStudy (anonymous):

what are you not understanding?

OpenStudy (anonymous):

and x too right ?

OpenStudy (anonymous):

x(x−(2−5√))(x−(2+5√))

OpenStudy (anonymous):

you can do this with any equation with specific roots if you wanted an equation with roots of 5, and 3 (x-5)(x-3)

OpenStudy (anonymous):

if one of the roots is zero then yes

OpenStudy (anonymous):

if you had roots -5, 2 (x-(-5)) (x-2) = (x+5)(x-2)

OpenStudy (anonymous):

then I sub the point in to find a right?

OpenStudy (anonymous):

no foil it out first x^(3)-4x^(2)-x = y then sub x= 2 into it and see how far it is away from 20 you get -10 so take your original equation and mutliply the entire thing by something that makes it equal to 20, -2 will do that :) -2(x^(3)-4x^(2)-x)

OpenStudy (anonymous):

now you have your answer

OpenStudy (anonymous):

do you get it? it really isn't hard at all

OpenStudy (anonymous):

where did u get this : x^(3)-4x^(2)-x = y

OpenStudy (anonymous):

1. Put roots into the form (x - (root))(x - (root)) 2. Sub the x value of the point into the equation and see what you get as a y value 3. multiply the entire equation by something that will give you the y value you are looking for. Well if you expended the equation you just generated for this question you would know

OpenStudy (anonymous):

i think u right, sorry but still i don't understand it :(

OpenStudy (anonymous):

I dont think I can explain it anymore than I already have, if you dont understand the concept take a break and come back and read it

OpenStudy (anonymous):

how about this ? create a cubic polynomial inequality for which x=3 or x<=-2 is the solution.

OpenStudy (anonymous):

you can have two solutions for that (x - 3)(x - (-2))(x - 3) = (x -3)(x + 2)(x-3) or (x - 3)(x - (-2))(x - (-2)) = (x - 3)(x + 2)(x + 2) if you expanded this you would end up with a cubic root equation with the roots specified